# Almost Linear Time Consistent Mode Estimation and Quick Shift Clustering

### v1

@@ Abstract
Keywords:
Density Based Clustering Locality Sensitive Hashing Quick Shift Clustering Mode Estimation
@@ 1 Introduction
Density-based clustering algorithms are fundamental tools in data analysis due to their ability to identify clusters of arbitrary shapes. The most popular density-based clustering method is DBSCAN [8][18], which defines clusters based on the concept of "density-reachability." Mean Shift [3][4][13] is another density-based clustering algorithm that moves each point to the densest area in its vicinity, based on kernel density estimation which is computationally challenging due to its iterative nature and the need for density estimation, making it much less scalable than DBSCAN.
To overcome this issue, Quick Shift [19][14] generalizes Mean Shift by constructing a hierarchical clustering tree on density estimates. However, all these methods are computationally expensive, making them less scalable for large datasets.
Jiang [14] established the consistency of the Quick Shift algorithm. This result allows the consistency analysis for various density estimators, including k-nearest neighbor [5]. Esfandiari et al. [7] demonstrated the use of certain types of LSH for fast density estimation, enabling efficient density-based clustering. Jang and Jiang [12] introduced DBSCAN++, a modification of DBSCAN that computes densities for a subset of points, reducing computational cost while maintaining performance. Xu and Pham [20] proposed sDBSCAN, a scalable density-based clustering algorithm in high dimensions using random projections. These works highlight the importance of efficient density estimation for scalable clustering.
Building upon these, we utilize hashing-based kernel density estimators [2][1] to develop a fast and consistent mode estimator and extend the Quick Shift algorithm as a result, achieving almost linear time complexity while preserving consistency.
@@ 2.1 Locality Sensitive Hashing
Approximate Near Neighbor Search allows for efficient data structures with query time sublinear in $n$ and polynomial dependence on $d$. A classic technique for solving $(c,r)$-ANN is Locality-Sensitive Hashing (LSH), introduced by Indyk and Motwani [11]. The main idea behind LSH is to use random space partitions such that pairs of points within distance $r$ are more likely to be hashed to the same bucket than pairs of points at a distance greater than $cr$.
Proposition 1 (Optimal LSH for $(c,r)$-ANNS [17])
@@ 2.2 Kernel Density Estimation
Proposition 2 (Approximate KDE via LSH [1])
@@ 3 Algorithm
Providing a dataset $X\subset\mathbb{R}^{d}$, there exist an algorithm (Algorithm 1) that preforms density based clustering in time and space $O(dn^{1+o(1)})$.
@@ 4 Theoretical Analysis
These properties form the foundation for consistent clustering guarantees, analogous to those established in the exact KDE setting [14], but now achieved with sub-quadratic computational complexity by virtue of using the approximate KDE as an oracle.
@@
with probability at least $1-1/n$ for all $q\in\mathcal{X}$, provided that $h\geq(\log n/n)^{1/d}$.
Under Assumptions 13, standard uniform convergence arguments (see, e.g., [14]) imply that there exists a constant $C^{\prime}>0$ such that with probability at least $1-1/n$
@@ 4.1 Mode Estimation
Quick Shift (see, e.g., [19]) is an iterative procedure that assigns each sample $x_{i}$ to a nearby point in its $\tau$-ball with strictly higher density. In our algorithm, each density evaluation is computed via $\tilde{f}(x)$. We denote by $\hat{M}$ the set of estimated modes returned by the algorithm.
@@
Then there exists a constant $C>0$ such that with probability at least $1-1/n$ the Hausdorff distance between the true mode set $M$ and the estimated mode set $\hat{M}$ satisfies
@@
Let $\hat{f}_{h}(x)=\frac{1}{n\,h^{d}}\sum_{i=1}^{n}K\!\Bigl{(}\frac{x-x_{i}}{h}\Bigr{)}$ be the classical KDE. Under Assumptions 1 and 2, standard results (e.g., Theorem 1, [14]) guarantee that with probability at least $1-1/n$,
@@ 4.2 Assignment of Points to Modes
    $\sup_{x\in S+B(0,r)}f(x)\;\;<\;\;\infty_{x\in B(x_{1},r)\,\cup\,B(x_{2},r)}f(x)\;-\;\delta.$
Equivalently, one may view $S$ as the region forming a “deep valley” separating $x_{1}$ from $x_{2}$, whose density is at least $\delta$ below the density near $x_{1}$ or $x_{2}$, plus a margin of width $r$. The main technical claim is that if two points $x_{1}$ and $x_{2}$ are $(r_{s},\delta)$-separated, then with high probability there is no directed path in $(G,\tilde{f})$ from $x_{1}$ to $x_{2}$.
@@
Let $(G,\tilde{f})$ be the directed cluster tree constructed by Quick Shift using the approximate densities $\tilde{f}$ from Assumption 3. Under the same conditions as Theorem 4.1, there exists a constant $C>0$ such that with probability at least $1-1/n$, if $x_{1}$ and $x_{2}$ are $(r_{s},\delta)$-separated, there is no directed path from $x_{1}$ to $x_{2}$ in $G$, provided $\delta>C\,\epsilon\,\sup_{x\in\mathcal{X}}\hat{f}(x)$.
@@
$f(y_{j}^{*})\;\;\leq\;\;\sup_{x\,\in\,S+B(0,r_{s})}f(x)\;\;<\;\;\infty_{x\in B(x_{1},r_{s})\cup B(x_{2},r_{s})}f(x)\;-\;\delta.$
Thus, if $\|x_{1}-y_{j}^{*}\|\leq\|x_{1}-x_{2}\|+\|x_{2}-y_{j}^{*}\|$, then in fact such a $y_{j}^{*}$ is forced to be on any path that attempts to connect $x_{1}$ to $x_{2}$. Now, if we compare $\tilde{f}(y_{j}^{*})$ and $\tilde{f}(x_{1})$, by KDE uniform bounds [15],
@@
An immediate corollary is that points $(r_{s},\delta)$-separated from a given mode $x_{0}$ cannot join $x_{0}$’s tree in $G$. Indeed, if $x$ were assigned to $x_{0}$, there would exist a directed path $x\to\cdots\to x_{0}$ in $(G,\tilde{f})$, violating Theorem 4.2. Hence, each point is constrained to remain in the basin of attraction of exactly those modes not separated from it by a deep valley.
@@
In the absence of approximation error (i.e. $\epsilon=0$), these results coincide exactly with the analysis in [14]. The only additional requirement here is that $\delta$ exceed a constant times $\epsilon$ (and the usual KDE deviation term), ensuring that approximate comparisons preserve the strict density gap. Thus, the same geometric intuition that “deep and wide valleys” prevent two points from being assigned to the same root remains valid under approximate density evaluations.
@@ 5 Experiments
We evaluate the proposed algorithm on two tasks: clustering, and image segmentation on benchmark data as the most well known tasks for Mean Shift variants. All experiments were conducted on a standard workstation (16 GB RAM, 2.4 GHz CPU).
@@ 5.1 Clustering
We compare the proposed algorithm against other popular density-based clustering algorithm and K-means as an scalable option on various clustering tasks in Table 1. These comparisons are made using the Scikit-Learn [^1] Scikit-Learn Webpage and Scikit-Image  [^2] Scikit-Image Webpage implementation of Quick Shift and our own implementation of LSH-Quickshift in C++ wrapped as a Python package using Pybind11, and used FAISS library [6] for LSH implemention. To measure the quality of a clustering result, we use the Adjusted Rand Index (ARI) [10] and the Adjusted Mutual Information (AMI) [16] scores, comparing the clustering with the partitioning induced by the labels of the data points. The benchmark datasets we use are labeled datasets from UCI Repository  [^3] UCI Machine Learning Repository Webpage, and we only cluster the features.
@@ 5.2 Image segmentation
We compare the proposed method to a number of baselines for unsupervised image segmentation in Figure 1. We include Felzenszwalb [9], Quick Shift [19], and Mean Shift [3], three popular image segmentation procedures from the Python, Scikit-Image library. For image segmentation, we run each algorithm on a preprocessed image with each pixel represented in a 5D $(r,g,b,x,y)$ color channel and spatial coordinates space and at maximum the size of our images was 46,500 pixels from the Berkeley Segmentation Dataset Benchmark (BSDS500)  [^4] Berkeley Segmentation Dataset Webpage. For each algorithm, the returned clusters are taken as the segments. Our image segmentation experiments show that LSH-Quick Shift is able to produce segmentations that are nearly identical to that of Mean Shift.
@@ 6 Conclusion
One may combining LSH-QuickShift with sub-sampling techniques and random projections and further enhance scalability and robustness, particularly in extremely high-dimensional settings with large data sets, and exploring the algorithm’s performance within massive parallel computing frameworks could lead to significant improvements in processing large-scale datasets, addressing both time and space complexity challenges. Also, applying these methods to various data modalities beyond numerical and image data, such as text or time-series data, could demonstrate its versatility and effectiveness across different domains.

### v2

@@ 1 Introduction
Density-based clustering algorithms are fundamental tools in data analysis due to their ability to identify clusters of arbitrary shapes. The most popular density-based clustering method is DBSCAN [ester1996density][schubert2017dbscan], which defines clusters based on the concept of "density-reachability." Mean Shift [cheng1995mean][comaniciu2002mean][jang2021meanshift++] is another density-based clustering algorithm that moves each point to the densest area in its vicinity, based on kernel density estimation which is computationally challenging due to its iterative nature and the need for density estimation, making it much less scalable than DBSCAN.
To overcome this issue, Quick Shift [vedaldi2008quick][jiang2017consistency] generalizes Mean Shift by constructing a hierarchical clustering tree on density estimates. However, these methods are computationally expensive, making them less scalable for large datasets.
Jiang [jiang2017consistency] established the consistency of the Quick Shift algorithm. This result allows the consistency analysis for various density estimators, including k-nearest neighbor [dasgupta2014optimal]. Esfandiari et al. [esfandiari2021DBSCAN] demonstrated the use of certain types of LSH for fast density estimation, enabling efficient density-based clustering. Jang and Jiang [jang2019dbscan++] introduced DBSCAN++, a modification of DBSCAN that computes densities for a subset of points, reducing computational cost while maintaining performance. Xu and Pham [xu2024scalable] proposed sDBSCAN, a scalable density-based clustering algorithm in high dimensions using random projections. These works highlight the importance of efficient density estimation for scalable clustering.
Building upon these, we utilize hashing-based kernel density estimators [charikar2017hashing][charikar2020kernel] to develop a fast and consistent mode estimator and extend the Quick Shift algorithm as a result, achieving almost linear time complexity while preserving consistency.
@@ 2.1 Locality Sensitive Hashing
Approximate Near Neighbor Search allows for efficient data structures with query time sublinear in $n$ and polynomial dependence on $d$. A classic technique for solving $(c,r)$-ANN is Locality-Sensitive Hashing (LSH), introduced by Indyk and Motwani [indyk1998approximate]. The main idea behind LSH is to use random space partitions such that pairs of points within distance $r$ are more likely to be hashed to the same bucket than pairs of points at a distance greater than $cr$.
Proposition 1(Optimal LSH for $(c,r)$-ANNS [o2014optimal])
@@ 2.2 Kernel Density Estimation
Proposition 2(Approximate KDE via LSH [charikar2020kernel])
@@ 3 Algorithm
Providing a dataset $X\subset\mathbb{R}^{d}$, there exist an algorithm (Algorithm 1) that performs density based clustering in time and space $O(dn^{1+o(1)})$.
@@ 4 Theoretical Analysis
These properties form the foundation for consistent clustering guarantees, analogous to those established in the exact KDE setting [jiang2017consistency], but now achieved with sub-quadratic computational complexity by virtue of using the approximate KDE as an oracle.
@@
with probability at least 11/n1-1/n for all $q\in\mathcal{X}$, provided that $h\geq(\log n/n)^{1/d}$.
Under Assumptions 13, standard uniform convergence arguments (see, e.g., [jiang2017consistency]) implies that there exists a constant $C^{\prime}>0$ such that with probability at least 11/n1-1/n
@@ 4.1 Mode Estimation
Quick Shift (see, e.g., [vedaldi2008quick]) is an iterative procedure that assigns each sample $x_{i}$ to a nearby point in its $\tau$-ball with strictly higher density. In our algorithm, each density evaluation is computed via $\tilde{f}(x)$. We denote by $\hat{M}$ the set of estimated modes returned by the algorithm.
@@
Then there exists a constant $C>0$ such that with probability at least 11/n1-1/n the Hausdorff distance between the true mode set $M$ and the estimated mode set $\hat{M}$ satisfies
@@
Let $\hat{f}_{h}(x)=\frac{1}{n\,h^{d}}\sum_{i=1}^{n}K\!\Bigl(\frac{x-x_{i}}{h}\Bigr)$ be the classical KDE. Under Assumptions 1 and 2, standard results (e.g., Theorem 1, [jiang2017consistency]) guarantee that with probability at least 11/n1-1/n,
@@ 4.2 Assignment of Points to Modes
    $\sup_{x\in S+B(0,r)}f(x)\;\;<\;\;\inf_{x\in B(x_{1},r)\,\cup\,B(x_{2},r)}f(x)\;-\;\delta.$
[…]
@@
[…]
@@
$f(y_{j}^{*})\;\;\leq\;\;\sup_{x\,\in\,S+B(0,r_{s})}f(x)\;\;<\;\;\inf_{x\in B(x_{1},r_{s})\cup B(x_{2},r_{s})}f(x)\;-\;\delta.$
Thus, if $\|x_{1}-y_{j}^{*}\|\leq\|x_{1}-x_{2}\|+\|x_{2}-y_{j}^{*}\|$, then in fact such a $y_{j}^{*}$ is forced to be on any path that attempts to connect $x_{1}$ to $x_{2}$. Now, if we compare $\tilde{f}(y_{j}^{*})$ and $\tilde{f}(x_{1})$, by KDE uniform bounds [jiang2017uniform],
@@
[…]
@@
In the absence of approximation error (i.e. $\epsilon=0$), these results coincide exactly with the analysis in [jiang2017consistency]. The only additional requirement here is that $\delta$ exceed a constant times $\epsilon$ (and the usual KDE deviation term), ensuring that approximate comparisons preserve the strict density gap. Thus, the same geometric intuition that “deep and wide valleys” prevent two points from being assigned to the same root remains valid under approximate density evaluations.
@@ 5 Experiments
We evaluate the proposed algorithm on two tasks: clustering, and image segmentation on benchmark data as the most well-known tasks for Mean Shift variants. All experiments were conducted on a standard workstation (16 GB RAM, 2.4 GHz CPU).
@@ 5.1 Clustering
We compare the proposed algorithm against other popular density-based clustering algorithm and K-means as an scalable option on various clustering tasks in Table 1. These comparisons are made using the Scikit-Learn [^1] Scikit-Learn Webpage and Scikit-Image  [^2] Scikit-Image Webpage implementation of Quick Shift and our own implementation of LSH-Quickshift in C++ wrapped as a Python package using Pybind11, and used FAISS library [faiss] for LSH implemention. To measure the quality of a clustering result, we use the Adjusted Rand Index (ARI) [ARI] and the Adjusted Mutual Information (AMI) [AMI] scores, comparing the clustering with the partitioning induced by the labels of the data points. The benchmark datasets we use are labeled datasets from UCI Repository  [^3] UCI Machine Learning Repository Webpage, and we only cluster the features.
@@ 5.2 Image segmentation
We compare the proposed method to a number of baselines for unsupervised image segmentation in Figure 1. We include Felzenszwalb [felzenszwalb2004efficient], Quick Shift [vedaldi2008quick], and Mean Shift [cheng1995mean], three popular image segmentation procedures from the Python, Scikit-Image library. For image segmentation, we run each algorithm on a preprocessed image with each pixel represented in a 5D $(r,g,b,x,y)$ color channel and spatial coordinates space and at maximum the size of our images was 46,500 pixels from the Berkeley Segmentation Dataset Benchmark (BSDS500)  [^4] Berkeley Segmentation Dataset Webpage. For each algorithm, the returned clusters are taken as the segments. Our image segmentation experiments show that LSH-Quick Shift is able to produce segmentations that are nearly identical to that of Mean Shift.
@@ 6 Conclusion
One may combine LSH-QuickShift with sub-sampling techniques and random projections and further enhance scalability and robustness, particularly in extremely high-dimensional settings with large data sets, and exploring the algorithm’s performance within massive parallel computing frameworks could lead to significant improvements in processing large-scale datasets, addressing both time and space complexity challenges. Also, applying these methods to various data modalities beyond numerical and image data, such as text or time-series data, could demonstrate its versatility and effectiveness across different domains.